1. What is the relationship between the Ksp expressions for a chemical reaction and its reverse chemical reaction?



1
Expert's answer
2022-03-20T06:37:23-0400

For an equilibrium reaction

"aA+bB \longleftrightarrow cC+dD"


"K_{sp}= \frac{[p_c]^c[p_D]^d}{[p_A]^a[p_B]^b}"


for the reverse reaction we can write value of KSp as follows


"K^{'}_{sp}= \frac{[p_A]^a[p_B]^b}{[p_C]^c[p_D]^d}"


"K_{sp}= \frac{1}{K^{'}_{sp}}"